The average length of time for students to register for fall classes at a certain college

has been 50 minutes with a standard deviation of 10 minutes. A new registration

procedure using modern computing machines is being tried. If a random sample of

12 students had an average registration time of 42 minutes with a standard

deviation of 11.9 minutes under the new system, test the hypothesis that the

population mean is now less than 50, using a level of significance of (a) 0.05, and (b)

0.01. Assume the population of time to be normal.


1
Expert's answer
2021-06-14T16:22:31-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\geq50"

"H_1:\mu<50"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha=0.05," and the critical value for a left-tailed test for "\alpha=0.05" and "df=n-1=12-1=11" degrees of freedom is "t_c=-1.795885." The rejection region for this left-tailed test is "R=\{t: t<-1.795885\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{42-50}{11.9/\sqrt{12}}\approx-2.328808"

Since it is observed that "t=-2.328808<-1.795885=t_c," it is then concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the population mean "\mu" is less than 50, at the "\alpha=0.05" significance level.


Using the P-value approach: The p-value for left-tailed, "t=-2.328808, df=11," "\alpha=0.05" is "p=0.01997," and since "p=0.01997<0.05," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is less than 50, at the "\alpha=0.05" significance level.


Based on the information provided, the significance level is "\alpha=0.01," and the critical value for a left-tailed test for "\alpha=0.01" and "df=n-1=12-1=11" degrees of freedom is "t_c=-2.718078." The rejection region for this left-tailed test is "R=\{t: t<-2.718078\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{42-50}{11.9/\sqrt{12}}\approx-2.328808"

Since it is observed that "t=-2.328808>-2.718078=t_c," it is then concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that the population mean "\mu" is less than 50, at the "\alpha=0.01" significance level.


Using the P-value approach: The p-value for left-tailed, "t=-2.328808, df=11," "\alpha=0.01" is "p=0.01997," and since "p=0.01997>0.01," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu" is less than 50, at the "\alpha=0.01" significance level.